OpenStudy (anonymous):

Specify the set by description. {1, 9, 81, 729, 6561, ...} nonnegative powers of 9 powers of 9 nonnegative multiples of 9 counting numbers multiples of 9

OpenStudy (anonymous):

It looks like multiples of 9.

OpenStudy (anonymous):

that's what I got

OpenStudy (anonymous):

1 * 9 = 9 9 * 9 = 81 9 * 81 = 729

ganeshie8 (ganeshie8):

multiples of 9 : 0, 9, 18, 27, 36, 45,...

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Oh yeah, dur.

OpenStudy (anonymous):

yes, 0 is in there as well

OpenStudy (anonymous):

yeah 0 * 9

OpenStudy (anonymous):

I thought powers of 9 meant x^9

OpenStudy (anonymous):

I feel pretty stupid right now. half of my brain is like "powers of 9 you dimwit" and the other half is like "well wait just a mo"